OpenStudy (anonymous):

WILL FAN AND MEDAL A set of equations is given below: Equation H: y = −x + 2 Equation J: y = 3x − 4 Which of the following steps can be used to find the solution to the set of equations? a.−x = 3x − 4 b.−x +2 = 3x c.−x + 2 = 3x − 4 d.−x + 1 = 3x + 2

OpenStudy (welshfella):

Hint: if y = some expression in x and y also = another expression in x then these 2 expressions are equal to each other.

OpenStudy (anonymous):

I'm lost. What do you mean?

OpenStudy (welshfella):

a simple case:- if y = a and y = b then a = b because they are both equal to y

OpenStudy (anonymous):

Ok. But how does that explain this problem?

OpenStudy (welshfella):

well look at the question y = 2 different things right - the 2 equations have 2 different expressions of x equal to y. now check out the options to see if these 2 expressions have an '=' between them

OpenStudy (anonymous):

I guess, to be honest I'm just very confused about this whole system and was wondering if somebody could take me through the steps

OpenStudy (welshfella):

well there only one step - refer back to y = a and y = b instead a theres -x + 2 and instead of b there is what? and these are equal

OpenStudy (anonymous):

b would be 3x = -4?

OpenStudy (welshfella):

no 3x - 4

OpenStudy (anonymous):

Right, sorry that's what i meant

OpenStudy (welshfella):

ok so the next step is to put these equal

OpenStudy (welshfella):

look through the options to see which one does that

OpenStudy (anonymous):

So it would be C.?

OpenStudy (welshfella):

yes

OpenStudy (welshfella):

y = -x + 2 y = 3x - 4 therefor -x + 2 = 3x - 4

OpenStudy (anonymous):

Thank you!!!!!!!!!!

OpenStudy (welshfella):

get it?

OpenStudy (welshfella):

yw